# Scalable Lattice Sampling using Factorized Generative Models

### v3

@@ Abstract
Lattice models in Physics are mathematical models that contain information in the form of a lattice $\bm{\phi}$ characterized by Boltzmann distribution defined with the help of a Hamiltonian $H(\bm{\phi})$ or action $s(\bm{\phi})$.
$p(\bm{\phi};\bm{\theta})\propto e^{-H(\bm{\phi};\bm{\theta})}$ (1)
Each lattice site is a random variable that could be binary as in the Ising model, circular as in the XY model or real-valued as in scalar $\phi^{4}$ theory. The statistical properties of these lattices vary with parameters $\bm{\theta}$, showing drastic changes near certain regions of $\bm{\theta}$ called critical regions. Although these distributions are known only up to a normalizing constant, they can be sampled using statistical methods such as MCMC (Markov Chain Monte Carlo) [1, 2]. These methods provide convergence guarantees that the samples represent the target distribution $p(\bm{\phi};\bm{\theta})$. But these methods could be quite inefficient, especially in the critical regions. In the critical regions, the correlation between successive samples of the Markov chain (characterized by the autocorrelation time of the chain) becomes very large. This phenomenon is called critical slowing down [3, 4]. Certain algorithms, such as, Swendsen-Wang [5], Wolff [6], worm [7], loop [8], directed loop [9] and HMC (Hamiltonian Monte Carlo) [10], make global MCMC updates to tackle the problem of critical slowing down. However, these methods have not been addressed for large lattices.
Generative machine learning (ML) methods are emerging as promising tools to sample Boltzmann distributions [11]. ML methods, such as Gaussian mixture models (GMM) and normalizing flows (NF), providing exact model probabilities, can be used in Metropolis-Hastings (MH) algorithm to propose samples which are accepted or rejected. Such algorithms provide convergence guarantees too. NF-based methods have been used for $\phi^{4}$ theory[12, 13, 14, 15, 16, 17, 18, 19, 20] and Gauge theory[21, 22, 23, 24, 25, 26, 27, 28, 29] in lattice field theory, and for general Monte Carlo sampling [30]. Self-learning Monte Carlo [31], [32] and Restricted Boltzmann Machines [33] also aim at efficient sampling of lattices using ML methods.
On the other hand, there are ML approaches that do not provide exact model probability, and hence, can not be used in the MH algorithm. These approaches are prone to biases. These include likelihood-free methods, such as, generative adversarial networks [34, 35, 36, 26] and likelihood-based methods, such as, variational autoencoders (VAEs) [37] and Boltzmann generators [38]. These methods produce one-shot samples, as opposed to Markov chains.
To address critical slowing down, some ML methods [39] train the generative model for every $\bm{\theta}$. Since training samples are not available for every $\bm{\theta}$, Reverse Kullback Leibler divergence (RKLD) based learning is used to train the model for a desired $\bm{\theta}$. This approach could be unstable. As an alternative, the use of conditional generative models is a promising approach [34, 40, 41], where the model learns from samples in non-critical regions and extrapolates or interpolates to critical regions.
In other words, do these approaches work efficiently and efficaciously with the increase in lattice size and for different parameters $\bm{\theta}$ (such as quark masses) [42]? Larger lattices provide a finer spatial resolution, allowing one to examine short-distance behaviours more accurately. They also help mitigate finite-size effects, ensuring that simulation results remain reliable and closer to the infinite-volume limit. It facilitates the study of phenomena like confinement, chiral symmetry breaking, and the QCD phase transition, which require the exploration of large-scale structures.
The major hurdle for scalability of learning-based approaches seems to be the joint modeling of the entire lattice. We propose Parallelizable Block Metropolis-within-Gibbs (PBMG) algorithm that factorizes the joint distribution of the lattice and samples in a Metropolis-within-Gibbs setting, with proposals generated by an ML model conditioned on $\bm{\theta}$. Gibbs algorithm has previously been used with ML models to sample large graphs [43], albeit not for sampling from explicit distribution functions.
We present the PBMG method for sampling from conditional Boltzmann distributions over large lattices. We use conditional generative models, e.g., conditional GMMs and conditional NFs, to efficiently sample in critical regions. The conditional generative models are generally trained from given samples. However, to avoid generating ground truth samples for large lattices, we use RKLD-based training. The problem of mode collapse is avoided in PBMG because the local distributions are much simpler (uni-modal or very few modes) as compared to the joint distributions.
To validate the proposed approach, we apply it to 2-D lattices, namely, the XY model from statistical Physics and the scalar $\phi^{4}$ model from lattice field theory. In the XY model, each lattice site is a spin value (angle) that is locally dependent on the immediate neighbors. The lattice distribution is conditioned on temperature $T$ and exhibits phase transition with respect to $T$, where the magnetic susceptibility diverges. We apply PBMG with conditional rational quadratic spline (RQS) flows for XY lattices [44]. In the scalar $\phi^{4}$ theory, each lattice site is a real number locally dependent on the immediate neighbors. The lattice distribution is conditioned on parameters $\lambda$ and $m^{2}$. We study phase transition with respect to $\lambda$, where susceptibility diverges. We apply PBMG with conditional GMMs, conditioned on $\lambda$ and $m^{2}$, for $\phi^{4}$ theory.
The contributions of this work are multifaceted:
  • The introduction of the PBMG algorithm, offering a robust and efficient means to sample Boltzmann distributions over lattices.
  • The pioneering factorization of the joint distribution into local components, laying the foundation for scalable sampling in high-dimensional lattice systems.
  • Proposed an inference algorithm that samples from all the local distributions in a parallel fashion to propose the final lattice sample thereby cutting down significantly on the computational cost.
  • The strategic utilization of RKLD optimization to alleviate the need for ground truth samples.
@@ 1 Parallelizable Block Metropolis-within-Gibbs (PBMG)
Every site-kernel $K_{i}$ for each $i\in g$ and for every partition $g$ is a Metropolis-within-Gibbs kernel, which means that each site-kernel $K_{i}$ is a Gibbs kernel with a Metropolis-Hastings accept-reject step. Metropolis-within-Gibbs algorithms, also known as conditional Metropolis-Hastings, are well studied in the MCMC literature [45].
@@ 2.2 Modeling the Proposal distribution
We use Rational Quadratic Splines (RQS) as the transform $f$. An RQS flow is used to map a fixed interval onto another fixed interval. The intervals are partitioned into $K$ bins where each bin is characterized by a rational-quadratic function that increases monotonically. Furthermore, these functions are characterized by $K+1$ coordinates (referred to as knots) representing boundary values, $K-1$ derivative values $\mathbf{s}_{knots}$ at internal knots and the widths $\mathbf{w}_{bins}$ and the heights $\mathbf{h}_{bins}$ of the $K$ bins. For details on RQS flow, one can refer to [44]. The coefficients governing the RQS transformation’s behaviour are adaptable and learned through neural networks. The base distribution is chosen to be uniform, and the interval of the uniform distribution $w_{int}$ is also learnable. We use four different neural networks to learn these parameters, as shown in figure 2.
@@ 3 Application to the $\phi^{4}$ Theory
The lattice $\phi^{4}$ theory on a 2D lattice has undergone extensive investigation in both statistical mechanics and quantum field theory due to its intricate phase structure and critical behaviour. It demonstrates second-order phase transitions, characterized by alterations in system properties like magnetization or correlation length with varying coupling parameters. For a thorough comprehension of the lattice $\phi^{4}$ model, one can refer to the works [46, 47, 48]. In the following section, we will discuss the proposed ML-based sampling approach for the lattice $\phi^{4}$ theory.
@@ 4 Experiments and Results for PBMG-XY
Here, $C(k)$ is the auto-correlation function of the chain at lag $k$ and $C(0)$ is the auto-correlation at zero lag. Higher auto-correlation necessitates more samples to ensure accurate estimation of observables. After generating the ensemble, we carry out thinning i.e., select every $k$th sample, where $k$ depends on the amount of auto-correlation in the chain, resulting in a chain with an effective sample size (ESS) of 2000. Here, ESS = $N/(2\tau_{int})$. The relative precision and confidence level in estimation for an ESS of 2000 are 91.235% and 95% respectively according to calculations by Vats et al [49]. Here, burn-in is unnecessary as we initiate the Markov chain from a high probability region with all spins at zero.
@@
Table 1 shows the EMD and %OL values for the mean energy observable. The values reported are the mean values along with standard deviations of EMD and %OL across the 23 temperatures. We observe that the EMD between the ground truth and the model is quite low and the %OL between the two is quite high. We compare the results with the work done by Singh et. al [34] for the two lattice sizes that they conducted their experiments on, i.e., $L=8$ and $L=16$. We observe that the metric values produced by our model are better than the ImplicitGAN-T model proposed by Singh et al. All of this supports the high accuracy of our approach.
@@ 5.1 Experiment on Observables
We also calculate the Earth Mover Distance and %OL values between PBMG-$\phi^{4}$ and HMC for $\lambda=4.8$ and $m^{2}=-4$ for the observable absolute mean magnetization. The final chains used for calculations have an ESS of 2000. We use 40 bins for generating the histogram corresponding to each distribution. The results are given in Table 3. For PBMG-$\phi^{4}$ too, we observe that the EMD between the ground truth and the model is quite low and the %OL between the two is quite high. We also compare these metrics for $L=8$ with the RKLD-based NF approach proposed by Albergo et. al [50] and observe that PBMG-$\phi^{4}$ is either at par or better than their model.
@@ 6 Computational Cost Analysis
Many approaches [13, 39, 14, 21, 40, 41] either require re-training of proposed models for every new set of parameters, or train the proposed models for a particular lattice size and carry out inference on the same size. With our proposed approach, however, the same model, once trained, can be used for different parameters and lattice sizes. This leads to a significant reduction in computational costs. Below we discuss the computational costs during inference. The inference from the PBMG models is carried out on a low-end single GPU machine.

### v4

# Improvement of Heatbath Algorithm in LFT using Generative model

@@ Abstract
Lattice models in physics describe systems on a discrete grid, where configurations follow a Boltzmann distribution characterized by a Hamiltonian $H(\phi,\lambda)$ or action $s(\phi,\lambda)$. The statistical properties of these lattices vary with parameters $\lambda$, exhibiting significant changes near specific values of $\lambda$ known as critical regions, where phenomena like critical slowing down occur. In these regions, traditional sampling methods like MCMC, while providing convergence guarantees, become inefficient due to high autocorrelation times. Specialized algorithms, such as Swendsen-Wang, Wolff address this but have limitations, particularly for continuous symmetries. The Hamiltonian Monte Carlo (HMC) [neal_hmc], which makes global MCMC updates, is considered the state of the art for lattice QCD simulations. Recently, generative machine learning (ML) methods, including normalizing flows, have shown promise in efficiently sampling lattice field theories, with successful applications in $\phi^{4}$ [medvidovic2021generative, albergoflowbased, Albergo:2022qfi, Hackett:2021idh, Kim_Nicoli_Paper1, Kim_Nicoli_Paper2, Caselle:2023mvh, Reyes-Gonzalez:2023oei, Coccaro:2023vtb, japneet, Singha:2021nht, pawlowski2020reducing, Wang:2023ex] and gauge theory [Kanwar_2020, Eichhorn:2023uge, Abbott:2022zhs, Abbott:2022hkm, Abbott:2023thq, Wang:2023exq, Foreman:2021ljl, Foreman:2021rhs, Boyda:2022nmh].
The generative methods mentioned above aim to model the joint distribution of the entire lattice, which poses a significant scalability challenge for learning-based approaches. An alternative well known approach for sampling in local lattice models is the Heat Bath algorithm, which factorizes the lattice distribution, allowing each lattice site to be sampled based on its local neighboring sites. This approach is particularly efficient for discrete systems like the Ising model, where only two possible states exist. However, in continuous systems, a straightforward sampling method is unavailable, often requiring rejection-based methods. The difficulty of finding effective proposal distributions for continuous systems can lead to high rejection rates, increasing the simulation cost.
We introduce the Parallelizable Block Metropolis-within-Gibbs (PBMG) method for sampling local lattice models, which learns the conditional local distribution and can be applied to sample the entire lattice. This approach trains a conditional generative model on a simple one-dimensional distribution, greatly enhancing the efficiency of the learning process. By focusing on this simplified target distribution, we also mitigate the issue of mode collapse commonly encountered when sampling from the joint distribution. With properly trained model we can reduce the rejection rate seen in Heat Bath sampling. The PBMG operates within a Metropolis-within-Gibbs framework, using ML model-generated proposals conditioned on neibours sites as well as action parameters. The PBMG are conditional generative models, e.g., conditional GMMs and conditional NFs, we do not require any training samples from the target distribution.
To validate the proposed approach, we apply it to 2-D lattices, namely, the XY model from statistical Physics and the scalar $\phi^{4}$ model from lattice field theory.
@@ 1 Parallelizable Block Metropolis-within-Gibbs (PBMG)
In this section, we present the Heat Bath methods with clear notation, detailing both the target distribution and the ML model as the proposal.
@@
Every site-kernel $K_{i}$ for each $i\in g$ and for every partition $g$ is a Metropolis-within-Gibbs kernel, which means that each site-kernel $K_{i}$ is a Gibbs kernel with a Metropolis-Hastings accept-reject step. Metropolis-within-Gibbs algorithms, also known as conditional Metropolis-Hastings, are well studied in the MCMC literature [jones:2014].
@@ 2.2 Modeling the Proposal distribution
We use Rational Quadratic Splines (RQS) as the transform $f$. An RQS flow is used to map a fixed interval onto another fixed interval. The intervals are partitioned into $K$ bins where each bin is characterized by a rational-quadratic function that increases monotonically. Furthermore, these functions are characterized by $K+1$ coordinates (referred to as knots) representing boundary values, $K-1$ derivative values $\mathbf{s}_{knots}$ at internal knots and the widths $\mathbf{w}_{bins}$ and the heights $\mathbf{h}_{bins}$ of the $K$ bins. For details on RQS flow, one can refer to [neural_spline_flows]. The coefficients governing the RQS transformation’s behaviour are adaptable and learned through neural networks. The base distribution is chosen to be uniform, and the interval of the uniform distribution $w_{int}$ is also learnable. We use four different neural networks to learn these parameters, as shown in figure 2.
@@ 3 Application to the $\phi^{4}$ Theory
The lattice $\phi^{4}$ theory on a 2D lattice has undergone extensive investigation in both statistical mechanics and quantum field theory due to its intricate phase structure and critical behaviour. It demonstrates second-order phase transitions, characterized by alterations in system properties like magnetization or correlation length with varying coupling parameters. For a thorough comprehension of the lattice $\phi^{4}$ model, one can refer to the works [AKDe, AKDe2, Vierhaus2010Simulation]. In the following section, we will discuss the proposed ML-based sampling approach for the lattice $\phi^{4}$ theory.
@@ 4 Experiments and Results for PBMG-XY
Here, $C(k)$ is the auto-correlation function of the chain at lag $k$ and $C(0)$ is the auto-correlation at zero lag. Higher auto-correlation necessitates more samples to ensure accurate estimation of observables. After generating the ensemble, we carry out thinning i.e., select every $k$th sample, where $k$ depends on the amount of auto-correlation in the chain, resulting in a chain with an effective sample size (ESS) of 2000. Here, ESS = $N/(2\tau_{int})$. The relative precision and confidence level in estimation for an ESS of 2000 are 91.235% and 95% respectively according to calculations by Vats et al [vats_et_al]. Here, burn-in is unnecessary as we initiate the Markov chain from a high probability region with all spins at zero.
@@
Table 1 shows the EMD and %OL values for the mean energy observable. The values reported are the mean values along with standard deviations of EMD and %OL across the 23 temperatures. We observe that the EMD between the ground truth and the model is quite low and the %OL between the two is quite high. We compare the results with the work done by Singh et. al [japneet] for the two lattice sizes that they conducted their experiments on, i.e., $L=8$ and $L=16$. We observe that the metric values produced by our model are better than the ImplicitGAN-T model proposed by Singh et al. All of this supports the high accuracy of our approach.
@@ 5.1 Experiment on Observables
We also calculate the Earth Mover Distance and %OL values between PBMG-$\phi^{4}$ and HMC for $\lambda=4.8$ and $m^{2}=-4$ for the observable absolute mean magnetization. The final chains used for calculations have an ESS of 2000. We use 40 bins for generating the histogram corresponding to each distribution. The results are given in Table 3. For PBMG-$\phi^{4}$ too, we observe that the EMD between the ground truth and the model is quite low and the %OL between the two is quite high. We also compare these metrics for $L=8$ with the RKLD-based NF approach proposed by Albergo et. al [albergo2021introduction] and observe that PBMG-$\phi^{4}$ is either at par or better than their model.
@@ 6 Computational Cost Analysis
Many approaches [albergoflowbased, Albergo, Albergo:2022qfi, Kanwar_2020, Singha:2023cql, Singha:2023xxq] either require re-training of proposed models for every new set of parameters, or train the proposed models for a particular lattice size and carry out inference on the same size. With our proposed approach, however, the same model, once trained, can be used for different parameters and lattice sizes. This leads to a significant reduction in computational costs. Below we discuss the computational costs during inference. The inference from the PBMG models is carried out on a low-end single GPU machine.